# Enhancing Diversity in Parallel Agents: A Maximum State Entropy Exploration Story

### v1

@@ 1 Introduction
A common formulation of the latter pure exploration setting is state entropy maximization (Hazan et al., 2019), in which the agent aims to maximize the entropy of the data distribution induced by its policy. The state entropy objective has been extensively considered in the literature, e.g., to provide data collection strategy for offline RL (Yarats et al., 2022; Park et al., 2024), experimental design (Tarbouriech & Lazaric, 2019), or transition model estimation (Tarbouriech et al., 2020), and as a surrogate loss for policy pre-training in reward-free settings (Mutti & Restelli, 2020a). All of the above uphold the status of state entropy maximization as a flexible tool for exploration. To this day, however, an important gap remains on how to deploy state entropy maximization on parallel simulators.
@@
Let us picture an illustrative example to guide our thoughts: A robotic arm control task. We may want the robot to be able to perform a wide range of tasks, such as manipulating different objects. A good strategy is to exploit a simulator of the robotic arm to learn good manipulation policies in advance, instead of relying on costly online interactions, in which exploration may be unsafe. Since not all of the future tasks may be known in advance, one could opt for state entropy maximization, in order to collect exploration data once-for-all. Collecting those data on complex simulators may be inefficient though, resulting in long simulation times. However, one crucial advantage of sampling in simulation rather than the physical system is that the sampling process does not have to be sequential.
@@
In many modern simulators (NVIDIA Isaac Sim, ; Genesis, 2024; Todorov et al., 2012), the process of learning is accelerated by instantiating vectorized instances of agents and environments simultaneously. Parallelization was shown to be a game changer approach that allowed for unprecedented performance in standard RL (Espeholt et al., 2018), and the use of a multitude of identical agents operating on copies of the same environment allowed RL systems to tackle increasingly complex and large-scale problems. While the maximum entropy formulation has been widely adopted in prior works, how to better take advantage of parallel exploration agents is still to be investigated.
@@
This is far from being a trivial problem.  Zhong et al. (2023) recently showed that not explicitly caring for redundant behaviors among the agents often results in inefficient use of computational resources, which is why parallelization is employed in the first place. Additionally, as highlighted in Rudin et al. (2022), solving complex problems necessitates intelligent strategies for constructing experience that accelerate algorithmic convergence. This challenge is further compounded in scenarios where access to real-world process data is limited, examples are scarce, or simulators are computationally expensive. Following these considerations, our study is centered around the questions below:
@@
How should parallel agents optimize their policies for efficient exploration? What is the best way to fully take advantage of parallelism for state entropy maximization?
@@
In this paper, we generally address this question by extending state entropy maximization to incorporate the role of parallel agents, enabling a deeper investigation into their exploration efficiency. First, we introduce a specific problem formulation that can explicitly leverage parallelism. Rather than considering agents independently, we look at them as a whole, and consider the distribution one virtual agent would induce if it was following a uniform mixture of the agents’ policies. Additionally, we theoretically characterize how the entropy of the distribution of state visits concentrates in the single-agent and parallel-agent cases and demonstrate the significant impact of parallelization on the rate of entropy stabilization and exploration diversity. Then, we move our attention on how to solve the problem of parallel state exploration in practice, by proposing a policy gradient strategy (Williams, 1992a) that explicitly optimizes for the objective at hand. Finally, we use illustrative experiments to show that the proposed algorithm is indeed able to optimize for the state entropy and that the induced behavior enables agents to balance individual exploration with inter-agent diversity. This synergy is particularly relevant when integrated with batch RL techniques (Levine et al., 2020).
Original Contributions.  Throughout the paper, we make the following contributions:
@@
  • We provide high-probability concentration bounds to describe the rate at which the entropy of empirical distributions converges to the entropy of their stationary distributions, in single and parallel agents settings, respectively. The proofs employ novel techniques that may be of independent interest (Section 4);
@@ 2 Preliminaries
We denote $N]:=\{1,2,\ldots,N\}$ for a constant $N<\infty$. We denote a set with a calligraphic letter $\mathcal{A}$ and its size as $|\mathcal{A}|$. We denote $\mathcal{A}^{T}:=\times_{t=1}^{T}\mathcal{A}$ the $T$-fold Cartesian product of $\mathcal{A}$. The simplex on $\mathcal{A}$ is denoted as $\Delta_{\mathcal{A}}:=\{p\in[0,1]^{|\mathcal{A}|}|\sum_{a\in\mathcal{A}}p(a)=1\}$ and $\Delta_{\mathcal{A}}^{\mathcal{B}}$ denotes the set of conditional distributions $p:\mathcal{A}\to\Delta_{\mathcal{B}}$. Let $X$ a random variable on the set of outcomes $\mathcal{X}$ and corresponding probability measure $p_{X}$, we denote as $\mathcal{H}(p_{X})=-\sum_{x\in\mathcal{X}}p_{X}(x)\log(p_{X}(x))$ the Shannon entropy of $X$.
@@
${d_{n}}(s)=\frac{1}{nT}\sum_{j=1}^{n}\sum_{t=0}^{T-1}\mathds{1}(S_{t,j}=s)\in\Delta_{\mathcal{S}},$
@@
In standard RL (Sutton & Barto, 2018), an agent interacts with an environment to maximize the (cumulative) reward collected from the MDP. In the absence of a reward, previous works (Hazan et al., 2019; Mutti & Restelli, 2020b; Mutti et al., 2021) investigated the effects of optimizing a different metric, namely the state entropy $\mathcal{H}_{\text{s}}$ defined as:
@@
The Jensen inequality relates the two formulations as $\mathcal{H}_{s}(\pi)\geq\mathcal{J}_{s}(\pi)$.
@@ 3 Problem Formulation
In this section, we extend the concept of state entropy to a novel objective that allows parallel agents to interact with independent copies of the environment. In order to do so, we first need to introduce the following objects:
@@
Collecting data from all agents induces an empirical (parallel) distribution over states which we denote as $d_{n,p}\in\Delta_{\mathcal{S}}$, so that it is possible to define the mixture distribution as the expectation of the empirical state distribution under the parallel policies, namely $d_{\pi_{p}}(s)=\mathbb{E}_{d_{n,p}\sim p_{\pi_{p}}}[d_{n,p}(s)]$. This objects allow us to define objectives specifically designed for PMDPs, that are then able to leverage their structural properties. First of all, we introduce the following:
@@
$\mathcal{H}_{s,p}(\pi)\triangleq\mathcal{H}(d_{\pi_{p}}):=-\sum_{s\in\mathcal{S}}d_{\pi_{p}}(s)\log d_{\pi_{p}}(s)$
@@
Let $\mathcal{M}_{m}$ a parallel MDP with an interaction over $n\in\mathbb{N}$ episodes, then the corresponding finite trials objective can be defined as:
@@
We are now ready to state the novel problem of interest, that we call Parallel State Entropy Maximization:
$\max_{\pi_{p}\in\Pi}\mathcal{J}_{p}(\pi_{p})=\max_{\pi_{p}\in\Pi}\left\{\mathbb{E}_{{d_{n,p}\sim p_{\pi_{p}}}}\mathcal{H}(d_{n,p})\right\}$ (1)
and we denote by $\pi_{p}^{\star}\in\arg\max_{\pi}\mathcal{J}_{p}(\pi)$ a parallel policy that maximizes the parallel states entropy. In the next sections, we will show that Parallel Learning Objectives enjoy extremely nice properties: first of all, they concentrate faster than their non-parallel counterparts, motivating the need for parallel exploration; moreover, they can be easily optimized in a decentralized fashion through policy gradient methods.
@@ 4 Fast Concentration of the Entropy in Parallel Settings
While the following theorems are instantiated for state distributions, they are of broader theoretical interest and apply equally to distributions over trajectories. Additionally, we adopt the convention of referring to the entropy and variance of distributions rather than those of random variables, as a slight abuse of terminology, since these quantities are technically defined for random variables but naturally extend to their associated distributions.
We start by analyzing the entropy of state distributions in MDPs, for which the following result holds (the proof can be found in Appendix A).
@@
Let $d_{\pi}$ be the (categorical) distribution induced by $\pi$ over the finite set $\mathcal{S}$ with $|\mathcal{S}|=S$, and let $d_{n}$ be the empirical distribution obtained from $n$ independent samples drawn from $d_{\pi}$.222nn represents the number of episodes for trajectory distributions or the number of steps for state distributions. Then, for any $\epsilon>0$, the following bound holds:
$\mathbb{P}\left(\mathcal{H}(d_{\pi})-\mathcal{H}(d_{n})>\epsilon\right)\leq 2S\exp\left(-n\cdot\frac{\epsilon^{2}\mathrm{Var}(d_{\pi})}{2S^{3}\mathcal{H}^{2}(d_{\pi})}\right),$
where $H(d_{n})$ and $H(d_{\pi})$ denote the entropy of the empirical and true distributions, respectively, and $\mathrm{Var}(d_{\pi})=\sum_{s\in[\mathcal{S}]}d_{\pi}(s)(1-d_{\pi}(s))$ is the variance of a random variable associated with the categorical distribution $d_{\pi}$. Furthermore, to ensure this concentration with confidence $1-\delta$, the number of samples $n$ must satisfy the following lower bound:
$n\geq\frac{2S^{3}\mathcal{H}^{2}(d_{\pi})}{\epsilon^{2}\mathrm{Var}(d_{\pi})}\cdot\ln\frac{2S}{\delta}.$
This theorem establishes an upper bound on the probability that the entropy difference between the true and empirical distributions exceeds $\epsilon$. Specifically, the probability of large deviations between these two entropies decreases exponentially with $n$, with the rate of convergence influenced by the entropy of the true distribution $d^{\pi}$. Notably, as $\lim_{\mathcal{H}(d_{\pi})\rightarrow 0}\frac{\mathcal{H}^{2}(d_{\pi})}{\mathrm{Var}(d_{\pi})}=0$, distributions with lower entropy require fewer samples for concentration, implying they are easier to approximate empirically.
@@
where the first term is the average entropy of the individual agent distributions, and the second term captures diversity among these distributions via the Kullback-Leibler divergence. By analyzing this decomposition, we can get insights into the potential advantages of parallel exploration versus single-agent exploration. To achieve high-entropy mixture distributions, we have two key strategies:
@@ 5 Policy Gradient for Parallel Exploration with a Handful of Trials
As stated before, a core motivation of this work is addressing the problem of exploration in practical scenarios, where a strong parallelization of the environment is used to overcome the difficulty of sampling from complex simulators or even physical instantiations of the agents. In such setting, it is crucial to achieve good performances over a finite number of realizations, and for this reason our attention now shifts towards finite trials objectives. Specifically, we will focus on the single trial case, i.e. $n=1$, and show that rather than instantiating $m$ copies of the same algorithm, the parallel learning objective allows for an alternative formulation, where agents explicitly leverage parallelization. The core idea lies in the implementation of a policy gradient of the parallel formulation of Eq. 3.3 on a set of parametric policies $\pi_{\theta_{i}}:\mathcal{S}\to\Delta_{\mathcal{A}}$, where $\theta_{i}\in\Theta_{i}\subset\mathbb{R}^{d}$ is a parameter vector in a d-dimensional space, uniquely associated with the agent $i$. The parallel policy is then defined as $\pi_{\theta}=(\pi_{\theta_{i}})_{i\in[m]}$.
In the following, we derive the gradient of the parallel objective with respect to the policy parameters. For the sake of simplifying the notation, we will refer to $d_{n,p}$ simply as $d_{p}$. Starting from the cost function:
@@
$\frac{\partial\mathcal{J}(\pi_{p})}{\partial\theta_{i}}=\frac{\partial}{\partial\theta_{i}}\int{p_{\pi_{p}}(d_{p})\mathcal{H}(d_{p})}$
@@
Then, after substituting back into original integral we get:
@@
Thus, the partial gradient update consists of an internal summation over the log of the $i$-th policy executed by agent $i$, as it interacts over the horizon $T$, experiencing various action-state combinations. This internal summation is further aggregated across the external summation over $K$ trajectories. This estimator approximates the true gradient by averaging the score function of sampled trajectories, weighted by the entropy of such induced state distribution. Limiting the parallel agents involved in the training loop to experience only one single trajectory each one, we mediate over a mini-batch $\mathcal{B}=\{(\tau_{i})\}_{i=1}^{n}$ iteration to reduce the variance of the reinforce estimator, with $B=|\mathcal{B}|$ its size. The pseudo-code of PGPSE (Policy Gradient for Parallel States Entropy Maximization) algorithm is reported in Algorithm 1.
Algorithm 1 PGPSE
[…]
$\pi_{\theta}\leftarrow\text{Softmax}(\theta)$
for $e\in\{1,\dots,N\}$ do
  for $itr\in\{1,\dots,B\}$ do
   for $k\in\{1,\dots,K\}$ do
    $\tau\sim\pi_{\theta}$  {Sample parallel trajectories}
    $\log\pi_{\theta_{i}}\leftarrow\sum_{t=1}^{T-1}\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid s_{t})$
    $d_{p}(s)\leftarrow\frac{1}{km}\sum_{j=1}^{m}\sum_{i=1}^{k}\sum_{t=1}^{T}\bm{1}(s_{t,i,j}=s)$
    $\nabla_{\theta}\mathcal{J}(\theta)\mathrel{+}=\log\pi_{\theta_{i}}\mathcal{H}(d_{p})$
   end for
  end for
  $\nabla_{\theta}\mathcal{J}(\theta)\leftarrow\frac{1}{B}\nabla_{\theta}\mathcal{J}(\theta)$
  $\theta\leftarrow\theta+\alpha\nabla_{\theta}\mathcal{J}(\theta)$
end for

### v2

@@ 1 Introduction
A common formulation of the latter pure exploration setting is state entropy maximization (Hazan et al., 2019), in which the agent aims to maximize the entropy of the data distribution induced by its policy. The state entropy objective has been widely studied in the literature, e.g., to provide data collection strategy for offline RL (Yarats et al., 2022; Park et al., 2023), experimental design (Tarbouriech & Lazaric, 2019), or transition model estimation (Tarbouriech et al., 2020), and as a surrogate loss for policy pre-training in reward-free settings (Mutti & Restelli, 2020). All of the above uphold the status of state entropy maximization as a flexible tool for exploration. To this day, however, an important gap remains in how to deploy state entropy maximization on parallel simulators.
@@
Let us picture an illustrative example to guide our thoughts: A robotic arm control task. We may want the robot to be able to perform a wide range of tasks, such as interacting with a diverse set of objects. A good strategy is to exploit a simulator of the robotic arm to learn good manipulation policies in advance, instead of relying on costly online interactions, in which exploration may be unsafe. Since not all of the future tasks may be known in advance, one could opt for state entropy maximization, in order to collect exploration data once and for all. Collecting those data on complex simulators may be inefficient though, resulting in long simulation times. However, one crucial advantage of sampling in simulation rather than the physical system is that the sampling process does not have to be sequential.
@@
In many modern simulators (NVIDIA Isaac Sim, ; Genesis, 2024; Todorov et al., 2012), the process of learning is accelerated by instantiating vectorized instances of agents and environments simultaneously. Parallelization was shown to be a game-changing approach that allowed for unprecedented performance in standard RL (Espeholt et al., 2018), and the use of a multitude of identical agents operating on copies of the same environment allowed RL systems to tackle increasingly complex and large-scale problems. While the maximum entropy formulation has been widely adopted in prior works, how to better take advantage of parallel exploration agents is still to be investigated.
@@
This is far from being a trivial problem.  Zhong et al. (2024) recently showed that not explicitly caring for redundant behaviors among the agents often results in inefficient use of computational resources, which is why parallelization is employed in the first place. Additionally, as highlighted in Rudin et al. (2022), solving complex problems necessitates intelligent strategies for constructing experience that accelerate algorithmic convergence. This challenge is further compounded in scenarios where access to real-world process data is limited, examples are scarce, or simulators are computationally expensive. Following these considerations, our study is centered around the following questions:
@@
How should parallel agents optimize their policies for efficient exploration? What is the best way to leverage parallelism in state entropy maximization?
@@
In this paper, we generally address this question by extending state entropy maximization to incorporate the role of parallel agents, enabling a deeper investigation into their exploration efficiency. First, we introduce a specific problem formulation that can explicitly leverage parallelism. Rather than considering agents independently, we look at them as a whole, and consider the distribution a single one virtual agent would induce if it was following a uniform mixture of the agents’ policies. Additionally, we theoretically characterize how the entropy of the distribution of state visits concentrates in the single-agent and parallel-agent cases and demonstrate the significant impact of parallelization on the rate of entropy stabilization and exploration diversity. Then, we shift our attention to how to solve the problem of parallel state exploration in practice, by proposing a policy gradient strategy (Williams, 1992) that explicitly optimizes for the objective at hand. Finally, we use illustrative experiments to show that the proposed algorithm is indeed able to optimize for the state entropy and that the induced behavior enables agents to balance individual exploration with inter-agent diversity. This synergy is particularly relevant when integrated with batch RL techniques (Levine et al., 2020).
Original Contributions.
Throughout the paper, we make the following contributions:
@@
  • We provide high-probability concentration bounds describing the rate at which the entropy of empirical distributions converges to the entropy of their stationary distributions. The proofs employ novel techniques that may be of independent interest (Section 4);
@@ 2 Preliminaries
We denote $N]:=\{1,2,\ldots,N\}$ for a constant $N<\infty$. We denote a set with a calligraphic letter $\mathcal{A}$ and its size as $|\mathcal{A}|$. We denote $\mathcal{A}^{T}:=\times_{t=1}^{T}\mathcal{A}$ the $T$-fold Cartesian product of $\mathcal{A}$. The simplex on $\mathcal{A}$ is denoted as $\Delta_{\mathcal{A}}:=\{p\in[0,1]^{|\mathcal{A}|}|\sum_{a\in\mathcal{A}}p(a)=1\}$ and $\Delta_{\mathcal{A}}^{\mathcal{B}}$ denotes the set of conditional distributions $p:\mathcal{A}\to\Delta_{\mathcal{B}}$. Let $X,X^{\prime}$ random variables on the set of outcomes $\mathcal{X}$ and corresponding probability measures $p_{X},p_{X^{\prime}}$, we denote the Shannon entropy of $X$ as $\mathcal{H}(X)=-\sum_{x\in\mathcal{X}}p_{X}(x)\log(p_{X}(x))$ and the Kullback-Leibler (KL) divergence as $D_{KL}(p_{X}\|p_{X^{\prime}})=\sum_{x\in\mathcal{X}}p_{X}(x)\log(p_{X}(x)/p_{X^{\prime}}(x))$.
@@
${d_{n}}(s)=\frac{1}{nT}\sum_{j=1}^{n}\sum_{t=0}^{T-1}\mathds{1}(S_{t,j}=s)\in\Delta_{\mathcal{S}}.$
@@
In standard RL (Sutton & Barto, 2018), an agent interacts with an environment to maximize the (cumulative) reward collected from the MDP. In the absence of a reward, previous works (Hazan et al., 2019; Mutti & Restelli, 2020; Mutti et al., 2021) investigated the effects of optimizing a different metric, namely the state entropy $\mathcal{H}_{\text{s}}$ defined as:
@@
The Jensen’s inequality relates the two formulations as $\mathcal{H}_{s}(\pi)\geq\mathcal{J}_{s}(\pi)$.
@@ 3 Problem Formulation
In this section, we extend the concept of state entropy to a novel objective able to incorporate parallel agents interacting with independent copies of the environment. In order to do so, we first need to introduce the following objects:
@@
Collecting data from all agents induces an empirical (parallel) distribution over states which we denote as $d_{n,p}\in\Delta_{\mathcal{S}}$, so that it is possible to define the mixture distribution as the expectation of the empirical state distribution under the parallel policies, namely $d_{\pi_{p}}(s)=\mathbb{E}_{d_{n,p}\sim p_{\pi_{p}}}[d_{n,p}(s)]$. This object allows us to define objectives specifically designed for PMDPs, that are then able to leverage their structural properties. First of all, we introduce the following:
@@
$\mathcal{H}_{s,p}(\pi)\triangleq\mathcal{H}(d_{\pi_{p}}):=-\sum_{s\in\mathcal{S}}d_{\pi_{p}}(s)\log d_{\pi_{p}}(s).$
@@
Let $\mathcal{M}_{m}$ a PMDP with an interaction over $n\in\mathbb{N}$ episodes, then the corresponding finite trials objective can be defined as:
@@
We are now ready to state the novel problem of interest:
Parallel State Entropy Maximization $\max_{\pi_{p}\in\Pi}\mathcal{J}_{p}(\pi_{p})=\max_{\pi_{p}\in\Pi}\left\{\mathbb{E}_{{d_{n,p}\sim p_{\pi_{p}}}}\mathcal{H}(d_{n,p})\right\}$ (1)
We denote by $\pi_{p}^{\star}\in\arg\max_{\pi}\mathcal{J}_{p}(\pi)$ a parallel policy that maximizes the parallel states entropy. In the next sections, we will show that Parallel Learning Objectives enjoy extremely nice properties: first of all, they concentrate faster than their non-parallel counterparts, motivating the need for parallel exploration; moreover, they can be easily optimized in a decentralized fashion through policy gradient methods.
@@ 4 Fast Concentration of the Entropy in Parallel Settings
While the following theorems are instantiated for state distributions, they are of broader theoretical interest and apply equally to distributions over trajectories. [^2] The number of samples $n$ will represent the number of episodes for trajectory distributions or the number of steps for state distributions. Additionally, we adopt the convention of referring to the entropy and variance of distributions rather than those of random variables, as a slight abuse of terminology, since these quantities are technically defined for random variables but naturally extend to their associated distributions.
We start by analyzing the entropy of state distributions in MDPs, for which the following result holds:
@@
Let $d_{\pi}$ be the (categorical) distribution induced by $\pi$ over the finite set $\mathcal{S}$ with $|\mathcal{S}|=S$, and let $d_{n}$ be the empirical distribution obtained from $n$ independent samples drawn from $d_{\pi}$. Then, for any $\epsilon>0$, the following bound holds: $\mathbb{P}\left(\mathcal{H}(d_{\pi})\!-\!\mathcal{H}(d_{n})\!>\!\epsilon\right)\leq 2S\exp\!\left(\!-n\frac{\epsilon^{2}\mathrm{Var}(d_{\pi})}{2S^{3}\mathcal{H}^{2}(d_{\pi})}\right)\!,$ where $H(d_{n})$ and $H(d_{\pi})$ denote the entropy of the empirical and true distributions, respectively, and $\mathrm{Var}(d_{\pi})=\sum_{s\in[\mathcal{S}]}d_{\pi}(s)(1-d_{\pi}(s))$ is the variance of a random variable associated with the categorical distribution $d_{\pi}$. Furthermore, to ensure this concentration with confidence $1-\delta$, the number of samples $n$ must satisfy the following lower bound: $n\geq\frac{2S^{3}\mathcal{H}^{2}(d_{\pi})}{\epsilon^{2}\mathrm{Var}(d_{\pi})}\cdot\ln\frac{2S}{\delta}.$
This theorem establishes an upper bound on the probability that the entropy difference between the true and empirical distributions exceeds $\epsilon$. Specifically, the probability of large deviations between these two entropies decreases exponentially with $n$, with the rate of convergence influenced by the entropy of the true distribution $d^{\pi}$. Notably, as $\lim_{\mathcal{H}(d_{\pi})\rightarrow 0}\frac{\mathcal{H}^{2}(d_{\pi})}{\mathrm{Var}(d_{\pi})}=0$, distributions with lower entropy require fewer samples for concentration, implying they are easier to approximate empirically. The full proof can be found in Appendix A.
@@
where the first term is the average entropy of the individual agent distributions, and the second term captures diversity among these distributions via the KL divergence. By analyzing this decomposition, we can get insights into the potential advantages of parallel exploration versus single-agent exploration. To achieve high-entropy mixture distributions, we have two key strategies:
@@ 5 Policy Gradient for Parallel Exploration with a Handful of Trials
As stated before, a core motivation of this work is addressing the problem of exploration in practical scenarios, where a strong parallelization of the environment is used to overcome the difficulty to sample from complex simulators or even physical instantiations of the agents. In such setting, it is crucial to achieve good performances over a finite number of realizations, and for this reason our attention now shifts towards finite trials objectives. Specifically, we will focus on the single trial case, i.e., $n=1$, and show that rather than instantiating $m$ copies of the same algorithm, the parallel learning objective allows for an alternative formulation, where agents explicitly leverage parallelization. The core idea lies in the implementation of a policy gradient of the parallel formulation of Eq. 3.3 on a set of parametric policies $\pi_{\theta_{i}}:\mathcal{S}\to\Delta_{\mathcal{A}}$, where $\theta_{i}\in\Theta_{i}\subset\mathbb{R}^{d}$ is a parameter vector in a d-dimensional space, uniquely associated with the agent $i$. The parallel policy is then defined as $\pi_{\theta}=(\pi_{\theta_{i}})_{i\in[m]}$.
In the following, we derive the gradient of the parallel objective with respect to the policy parameters. For the sake of simplifying the notation, we will refer to $d_{n,p}$ as $d_{p}$ for simplicity. Starting from the cost function:
@@
$\frac{\partial\mathcal{J}(\pi_{p})}{\partial\theta_{i}}=\frac{\partial}{\partial\theta_{i}}\int{p_{\pi_{p}}(d_{p})\mathcal{H}(d_{p})}.$
@@
Then, after substituting back into the original integral we get:
@@
Algorithm 1: Policy Gradient for Parallel States Entropy maximization (PGPSE) 1:Input: Episodes N, Trajectories K, Batch Size B, Learning Rate $\alpha$, Parameters $\theta=(\theta^{i})_{i\in[m]}$ 2:for $e\in\{1,\dots,N\}$ do 3:  for $itr\in\{1,\dots,B\}$ do 4:   for $k\in\{1,\dots,K\}$ do 5:    $\tau\sim\pi_{\theta}$  {Sample parallel trajectories} 6:    $\log\pi_{\theta_{i}}\leftarrow\sum_{t=1}^{T-1}\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid s_{t})$ 7:    $d_{p}(s)\leftarrow\frac{1}{km}\sum_{j,i,t=1}^{m,k,T}\bm{1}(s_{t,i,j}=s)$ 8:    $\nabla_{\theta}\mathcal{J}(\theta)\mathrel{+}=\log\pi_{\theta_{i}}\cdot\mathcal{H}(d_{p})$ 9:   end for 10:  end for 11:  $\nabla_{\theta}\mathcal{J}(\theta)\leftarrow\frac{1}{B}\nabla_{\theta}\mathcal{J}(\theta)$ 12:  $\theta\leftarrow\theta+\alpha\nabla_{\theta}\mathcal{J}(\theta)$ 13:end for 14:Output: Policies $\pi_{\theta}=(\pi^{i}_{\theta^{i}})_{i\in[m]}$
Thus, the partial gradient update consists of an internal summation over the log of the $i$-th policy executed by agent $i$, as it interacts over the horizon $T$, experiencing various action-state combinations. This internal summation is further aggregated across the external summation over $K$ trajectories. This estimator approximates the true gradient by averaging the score function of sampled trajectories, weighted by the entropy of such induced state distribution. By limiting parallel agents to experience only one trajectory per each, we average over a mini-batch $\mathcal{B}=\{(\tau_{i})\}_{i=1}^{n}$ to reduce the variance of the reinforce estimator, with $B=|\mathcal{B}|$ being the batch-size. The pseudo-code of the resulting Algorithm, called Policy Gradient for Parallel States Entropy maximization (PGPSE), is reported in Algorithm 5.